### v2

# Towards White Box Deep Learning

Abstract
This paper introduces semantic features as a candidate conceptual framework for building inherently interpretable neural networks. A proof of concept model for informative subproblem of MNIST consists of 4 such layers with the total of ~5K learnable parameters. The model is well-motivated, inherently interpretable, requires little hyperparameter tuning and achieves human-level adversarial test accuracy - with no form of adversarial training! These results and the general nature of the approach warrant further research on semantic features. The code is available at https://github.com/314-Foundation/white-box-nn.
1 Introduction
Machine learning systems used in safety-critical applications such as healthcare, military, autonomous vehicles or cybersecurity need to be reliable, i.e. they cannot fail unpredictably. The black box nature of state of the art (SOTA) deep neural networks (DNN) make it exceedingly hard to give such guarantees, which limits their usability to non-critical domains. The recent research in Explainable AI (XAI) focuses predominantly on post-hoc attribution methods which may provide explanations but, due to the complexity of SOTA, are hard to be converted to reliable fail-safes. This is a burning issue because of the wide susceptibility of DNNs to distribution shifts and adversarial attacks [18, 5] meaning that models can make arbitrarily incorrect predictions and their reported high test-time metrics don’t guarantee robust real-life performance111models often rely on spurious features that leak from test to train [6].. The recently popular Neurosymbolic AI [17] consists of higher-level methods that focus on augmenting neural network representations (machine perception) with background structured knowledge (reasoning) but don’t address the potential failures of the former.
Existing methods of improving DNN robustness are far from ideal. Adversarial Training [11] being one of the strongest empirical defenses against attacks [1] is slow, data-hungry [15, 8] and results in robust generalisation gap [13, 21] as models overfit to the train-time adversaries. Thus, the biggest challenge of empirical approaches is reliable evaluation of robustness as more and more sophisticated attacks break the previously successful defenses [3, 4].
To address these challenges the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [9]. However the exact certification is NP-Complete [20] and therefore has serious limitations in practice as it’s limited to relatively small problems and small networks and/or to probabilistic approaches. Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation. The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/. for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
The highlighted issues are relevant for datasets as simple as MNIST [16] or even some of its binary subsets [19] - despite being "toy" and almost linearly separable problems, they still remain elusive in terms of robustness, rising further questions about the true generalisation capabilities of SOTA as no neural network model can be honestly considered reliable and adequate333in terms of internal complexity; SOTA models define inappropriately complicated functions. for this allegedly "simple" problem.
1.1 Proposed approach
This paper proposes an intermediate approach to DNN reliability and introduces a candidate theoretical framework for building white box neural networks. Building and analysing a white box DNN would be similar to writing and debugging a procedural program - it doesn’t have to come with theoretical certificates but needs to be easily understandable and internalizable by the programmer. A good software has to be designed well enough to account for reasonable failure scenarios, handle foreseeable exceptions and, should an unforseen bug happen, be amenable for fixes. If the design and inner mechanics of a DNN are sufficiently clear then working with such model wouldn’t be very different from maintaining a standard application.
Current SOTA models are large, unwieldy black boxes that require prohibitive amounts of resources; yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135K neurons! This scale seems easily maintainable with proper abstractions. After all, the core algorithms used today are mostly unchanged since ~30 years! What if there really exists a better way?
To find it we must start with the basics and take the agile approach. Focusing on the simplest relevant problem cuts off the obfuscating details and speeds up the experimentation phase. With that in mind the Minimum Viable Dataset (MVD) for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [19] and therefore might be considered harder than the entire MNIST in terms of average adversarial accuracy444the preliminary results on the full MNIST are discussed in Appendix C..
1.2 The core idea
This paper introduces semantic features as a general idea for sharing weights inside a neural network layer. The rigorous definition will be given in the next section. The concept is similar to that of "inverse rendering" in Capsule Networks [14] where features have many possible states and the best-matching state has to be found. Identifying different states by the subsequent layers gives rise to controlled dimensionality reduction. Thus semantic features aim to capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time. This is in fact a pretty strong regularization. As shown in this paper, choosing appropriate layers of semantic features for the MVD results in what can be considered as a white box neural network.
The remainder of the paper is structured as follows:
  • Section 2. gives a rigorous definition of semantic feature and argues for the generality of the notion;
  • Section 3. builds a carefully motivated PoC 4-layer white box neural network for selected MVD;
  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and ablation;
  • Section 5. provides a short discussion of the results;
The paper ends with ideas for further research and acknowledgments.
2 Semantic features
In machine learning inputs are represented as tensors. Yet the standard Euclidean topology in tensor space usually fails to account for small domain-specific variations of features. For example shifting an image by 1 pixel is semantically negligible but usually results in a distant output in $L_{2}$ metric.
This observation inspires the following definition:
Definition:
A semantic feature $f_{\mathbf{P}\mathbf{L}}$ of dimension $s\in\mathbb{N}$ is a tuple $(f,\mathbf{P},\mathbf{L})$ where:
  • $f$ base, $f\in\mathbb{R}^{s}$,
  • $\mathbf{P}$ parameter set, $\mathbf{P}\subset\mathbb{R}^{r}$ for some $r\in\mathbb{N}$
  • $\mathbf{L}$ locality function : $\mathbf{P}\rightarrow\textrm{Aut}(\mathbb{R}^{s},\mathbb{R}^{s})$ where Aut is the family of differentiable automorphisms of $\mathbb{R}^{s}$
Additionally the feature $(f,\mathbf{P},\mathbf{L})$ is called differentiable iff $\mathbf{L}$ is differentiable (on some neighbourhood of $\mathbf{P}$).
Intuitively a semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [14]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias555by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
2.1 Examples of semantic features
To better understand the definition consider the following examples of semantic features:
  1. 1.
    real-valued $f_{\mathbf{P}\mathbf{L}}$
    • $f\in\mathbb{R}$
    • $\mathbf{P}$ a real interval $[p_{min},p_{max}]$ for $p_{min}<=0<=p_{max}$
    • $\mathbf{L}$ function that maps a real number $p$ to the function $g\mapsto g+p$ for $g\in\mathbb{R}$
  2. 2.
    convolutional $f_{\mathbf{P}\mathbf{L}}$
    • $f$ square 2D convolutional kernel of size $(c_{out},c_{in},k,k)$
    • $\mathbf{P}$ set of $2\times 2$ rotation matrices
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel666can be made differentiable as in [12, 7].
  3. 3.
    affine $f_{\mathbf{P}\mathbf{L}}$
    • $f$ vector representing a small 2D spatial feature (e.g. image patch containing a short line)
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices777preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images6
  4. 4.
    xor $f_{\mathbf{P}\mathbf{L}}$
    • $f$ single-entry vector of dimension $n\in\mathbb{N}$
    • $\mathbf{P}\subset\{0,1,\ldots,n-1\}$
    • $\mathbf{L}$ function that maps a natural number $p$ to the function that rolls elements of $g\in\mathbb{R}^{n}$ by $p$ coordinates to the right
  5. 5.
    logical $f_{\mathbf{P}\mathbf{L}}$
    • $f$ concatenation of $k$ single-entry vectors and some dense vectors in-between
    • $\mathbf{P}\subset\{0,1,\ldots,n_{0}-1\}\times\ldots\times\{0,1,\ldots,n_{k-1}-1\}$
    • $\mathbf{L}$ function that maps $p\in\mathbf{P}$ to the respective rolls of single-entry vectors (leaving the dense vectors intact)
2.2 Matching semantic features
The role of semantic features is to be matched against the input to uncover the general structure of the data. Suppose that we have already defined a function match: $(\mathbb{R}^{c},\mathbb{R}^{s})\rightarrow\mathbb{R}$ that measures an extent to which datapoint $d\in\mathbf{R}^{c}$ contains a feature $g\in\mathbb{R}^{s}$. Now let’s define
$\mathrm{SFmatch}(d,(f,\mathbf{P},\mathbf{L}))=\max_{p\in\mathbf{P}}\mathrm{match}(d,\mathbf{L}(p)(f))$ (1)
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates888i.e. $f_{p}$ is a conjunction of lower-level features.. Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks. This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
  • $f=[1,0,-1,-1]$;           $\mathbf{P}=\{0,1\}$;           $\mathbf{L}(0)(f)=[1,0,-1,-1]$ and $\mathbf{L}(1)(f)=[0,1,-1,-1]$.
It remains to characterise the match function. Usually it can be defined as the scalar product $d\cdot emb(g)$ where $emb:\mathbb{R}^{s}\rightarrow\mathbb{R}^{c}$ is some natural embedding of the feature space to the input space. In particular for our previous examples lets set the following:
  • for real-valued $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\mathrm{int}(d==g)$ where $d,g\in\mathbb{R}$
  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 999alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
  • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape
  • for logical $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot g$ for $g,d\in\mathbb{R}^{c}$
The occasional $L_{2}$ normalization in the above examples is to avoid favouring norm-expanding $\mathbf{L}(p)$.
3 Network architecture
In this section we will use semantic features to build a PoC white box neural network model that classifies MNIST images of "3" and "5" in a transparent and adversarially robust way. We will build the network layer by layer arguing carefully for every design choice. Note that the network construction is not an abstract process and must reflect the core characteristics of the chosen dataset. There is no free lunch, if we want to capture the semantics of the dataset we need to encode it in the architecture - more or less explicitly. The framework of semantic features allows to do this in a pretty natural way. Note that a single neural network layer will consist of many parallel semantic features of the same kind.
The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $5$K parameters.
3.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH101010the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
A real-valued $f_{\mathbf{P}\mathbf{L}}$ with match and $\mathbf{L}$ defined as in the previous section identifies numbers in certain interval around $f$. This means that a layer consisting of multiple real-valued $f_{\mathbf{P}\mathbf{L}}$, provided that the relevant intervals don’t overlap, can be represented as a single locally-constant real-valued function. This allows us to simplify the implementation of such layer and instead of learning the constrained parameters of multiple $f_{\mathbf{P}\mathbf{L}}$ we can learn a single parameterised real-valued function that serves as an entire layer. In order to group pixel intensities into 3 abstract values we can draw inspiration from the $\mathrm{softsign}(x)=\frac{x}{1+x}$ function or rather from it’s parameterised version $\mathrm{softsign}(x,t,s)=\frac{x-t}{s+x}$ and “glue” two of such step functions together to obtain (Parameterised) Two Step function shown in Figure 2. The exact implementation of the Two Step function is provided in Appendix A. We initialize the layer setting init_scales_div=10.
3.2 Convolutional Layer
A pixel-wise function is not enough to classify pixel as ON or OFF - the semantics of our MVD require a pixel to exist in a sufficiently bright region to be considered as ON. Therefore the next layer will consist of a single convolutional semantic feature $(f,\mathbf{P},\mathbf{L})$:
  • $f$ 2D kernel of shape $(2,1,5,5)$ (concatenation of $(g^{0},g^{1})$, both of shape $(1,1,5,5)$)
  • $\mathbf{P}$ fixed set of $k$ distinct rotations by a multiple of the $\frac{360^{\circ}}{k}$ angle, for $k=32$
  • $\mathbf{L}$ as defined earlier
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches111111a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
The layer is followed by ReLU activation. We initialize $g^{0}$ and $g^{1}$ as the identity kernel of shape $(1,1,5,5)$ plus Gaussian noise $\sim\mathcal{N}(0,0.1)$.
3.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small121212close to the identity augmented matrix affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
  • $f$ 2D image patch of shape $(1,20,20)$
  • $\mathbf{P}$ set of 32 affine $2\times 3$ augmented matrices
  • $\mathbf{L}$ as defined earlier
For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is zero-padded evenly on every side and identified with the resulting $28\times 28$ image. Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
The layer is followed by ReLU activation. We initialize affine matrices as $2\times 2$ identity matrix plus Gaussian noise $\sim\mathcal{N}(0,0.01)$ concatenated with $1\times 2$ matrix filled with Gaussian noise $\sim\mathcal{N}(0,1)$.
3.4 Logical Layer
After Affine Layer has extracted predictive shapes it remains to encode logical claims such as "number 3 is A or B and not C and not D" etc. Since the previous layer is learnable we can enforce a preferable structure on the input neurons to the Logical Layer. This layer will consist of 2 logical $f_{\mathbf{P}\mathbf{L}}$ (one for every label) defined as follows:
  • $f$ vector of dimension 8
  • $\mathbf{P}=\{0,1,2,3\}$
  • $\mathbf{L}$ as defined earlier
We initialize $f^{0}=[\frac{4}{10},0,0,0,-\frac{1}{10},-\frac{1}{10},-\frac{1}{10},-\frac{1}{10}]$ and $f^{1}=[-\frac{1}{10},-\frac{1}{10},-\frac{1}{10},-\frac{1}{10},\frac{4}{10},0,0,0,]$. The exact initialization values are less important then their sign and relative magnitude, however they should not be too large as we will train the network using CrossEntropyLoss.
4 Results
4.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability131313mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
Reliability curve
Learning curve
Figure 1: (left) Reliability test-time curve for the classic 40-step PGD Attack. (right) Learning curve.
4.2 Quantitative results
We don’t bother measuring the clean test metrics as it is a flawed approach that tells little about the true generalisation capability of the model. Every metric reported here is computed under strong adversarial regime.
The model achieves ~92% accuracy under AutoAttack [4] (with default parameters: norm=’inf’, eps=0.3, step_size=0.1). The metrics are almost identical for the much faster 40-step PGD Attack. Samples are usually misclassified with low confidence and therefore we compute a reliability curve - a precision-recall curve where the positive class is defined as the set of correctly classified samples (this means that the precision for 100% recall is exactly equal to the accuracy). It turns out that for 80% recall model achieves human-level 98% adversarial precision (Figure 1). This means that in practice human intervention would be required only for the easily detected 20% of adversarial samples to achieve human level reliability of the entire system (and most real-life samples are not adversarial).
4.3 Layer inspection
Lets inspect the trained model layers.
4.3.1 Two Step Layer
Figure A shows the initial and learned shapes of the Two Step function. The function smoothly thresholds input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON. Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
Figure 2: Two Step layer - initial and learned.
4.3.2 Convolutional Layer
Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
Figure 3: Learned convolutional kernels in 32 predefined poses.
4.3.3 Affine Layer
Figure 4 shows the learned affine base features. Figure B in the Appendix B shows learned features in 32 learned feature-specific poses. It’s hard to imagine substantially more meaningful features.
Figure 4: Learned affine base features.
4.3.4 Logical Layer
The two learned logical base features are as follows:
  • $f^{0}=[1.0700,0.0000,0.0000,0.0000,-0.7800,-0.7100,-0.8000,-0.8100]$
  • $f^{1}=[-0.8100,-0.7600,-0.8600,-0.8000,1.0500,0.0000,0.0000,0.0000]$
This is exactly as expected - $f_{\mathbf{P}\mathbf{L}}^{0}$ defines number "3" as an entity having large SFmatch with any of the first 4 affine semantic features and small SFmatch with all of the remaining 4 affine semantic features. For $f_{\mathbf{P}\mathbf{L}}^{1}$ its the other way round.
4.4 Decision boundary
Figure 5 shows the minimum perturbations required to change model’s predictions under Boundary Attack [2] for sample test images. The added perturbations are easily spotted and understood by humans. Note that by definition model is extremely unconfident on those samples and in practice the perturbations required to fool the entire system would have to be much larger.
Figure 5: Boundary Attack: minimal perturbations that flip model prediction (model confidence at around 50%).
Figure 6: Visualization of test data as "seen" by the first 3 consecutive layers.
Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and labels denote respectively:
  1. 1.
    adversarial input and predicted class probabilities
  2. 2.
    output of Two Step Layer
  3. 3.
    output of Convolutional Layer
  4. 4.
    top pose of affine features predicting label "3" (with the match value)141414these rows are actually overlaid with the middle row to make it easier to see what has been matched.
  5. 5.
    top pose of affine features predicting label "5" (with the match value)14
Note that in Figure 6 there are 3 misclassified examples and for every one of them it’s easy to understand why the model made the mistake.
4.5 Ablation study
Most of the heavy lifting regarding pixel-intensity alignment is done by the Two Step Layer which lifts adversarial robustness from 0% to ~86%. The Convolutional Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
Interestingly the number of affine semantic features can be reduced from 8 to 2 with the cost of just ~2%pt adversarial accuracy. Logical Layer can be replaced by a simple linear probe, however it needs a proper initialization to avoid the risk of dead neurons (which is high because of the exceedingly small amount of neurons).
The Gaussian noise augmentation is vital for the model to experience pixel intensity variations during training (boosts adversarial robustness by ~62%pt).
Despite being primarily designed to withstand the $L_{\infty}$ attack the model has considerable robustness against the standard $L_{2}$ AutoPGD attack (eps=1.5, eps_step=0.1) with ~89% adversarial accuracy.
5 Discussion
We’ve built a white box neural network that is naturally robust, reliable and inherently interpretable. Thanks to the notion of semantic features the construction of the model was as straightforward as writing a fully explainable procedural program and so are the inner workings of the network.
Affine Layer is the core component of the model and the main reason for its interpretability. However if applied directly to the input it is easily fooled by attacks. It’s white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 6 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature). Human eye appears to ignore such considerate changes in pixel intensities which means that the naive linear representation of pixel brightness is not semantically adequate. This is exactly how (and why) the first two layers of the model were designed - to fix the inherent vulnerability of otherwise perfectly interpretable affine features. These layers are themselves compatible with the framework of semantic features which suggests a large untapped potential of the notion.
The Logical Layer has been developed as an attempt to formalize the idea of an abstract hidden layer sparse activation structure and in that regard it is the most inspiring; however it might turn out to be too generic (on MNIST the properly-initialized linear probe seems to be sufficient).
6 Further research
Future work would have to involve more complex datasets and more nuanced semantic features. For example notice the analogy between affine semantic features and the SIFT [10] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional) are essentially performing input normalization. This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
There might be a concern that adding more layers to the network would reduce its interpretability. However the role of additional layers would be to widen the receptive field of the network. The Affine Layer defined earlier matches parts-at-locations of center-aligned objects. Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object). This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
Other ideas for further work include:
  • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction;
  • study non-affine perturbations of features, e.g. projective perturbations;
  • find adequate semantic topology for color space;
  • define semantic features for sound (inspiration for semantic invariants could be drawn from music theory);
  • define semantic features for text, e.g. capture semantic invariance of permutations of words;
  • study more complex logical semantic features in more elaborate scenarios;
  • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable;
7 Acknowledgments
This is work was done as an independent self-funded research.
However it needs to be noted that I had the opportunity to do preliminary investigation on adversarial robustness as a half-time activity during my employment at MIM Solutions. Special thanks go to Piotr Sankowski, who introduced me to the issue of adversarial vulnerability of neural networks, and both to Piotr Sankowski and Piotr Wygocki who - as my former employers - decided to trust me with that research as a half-time part of my contract. It was precisely thanks to this opportunity that I developed a strong intuition that modern neural network architectures are inherently flawed and need to be fundamentally re-examined.
I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk and members of the reddit community for the great early feedback on the first version of this paper.
Appendix A Implementation of Two Step Layer
class TwoStepFunction(Module):
"""Simplifiedreal-valuedSFLayerforinputsin[0,1]"""
def init_weights(self):
self.a0 = nn.Parameter(torch.tensor(0.5))
self.a1 = nn.Parameter(torch.tensor(0.5))
self.t0 = nn.Parameter(torch.tensor(0.25))
self.t1 = nn.Parameter(torch.tensor(0.75))
self.scales = nn.Parameter(
(torch.ones((4,)) / self.init_scales_div)
def softsign(self, x, threshold, scale, internal):
a = self.a1 if threshold > 0 else self.a0
x = x - threshold
x = x / (scale + x.abs())
if internal:
shift = threshold
shift = shift / (scale + shift.abs())
x = x + shift
x = a * x / shift.abs()
return x
div = x.abs().max() / (1 - a)
x = x / div
x = x + threshold.sgn() * a
return x
def transform(self, x):
return 2 * x - 1
def forward(self, x):
x = self.transform(x)
t0, t1 = self.transform(self.t0), self.transform(self.t1)
params = [
(t0, self.scales[0], False),
(t0, self.scales[1], True),
(t1, self.scales[2], True),
(t1, self.scales[3], False),
xs = [self.softsign(x, *p) for p in params]
masks = [
(x < t0).float(),
(t0 <= x).float() * (x < 0.0).float(),
(0.0 <= x).float() * (x < t1).float(),
(t1 <= x).float(),
x = sum(m * xx for (m, xx) in zip(masks, xs))
return x
Appendix B Learned Affine Features with learned poses
Appendix C Preliminary results for the entire MNIST
Let’s review the results for a straightforward adaptation of previously defined model for the full MNIST:
  • Affine Layer is expanded to 50 affine semantic features;
  • Logical Layer is replaced by a linear probe with 10 output classes;
  • Size of affine features is reduced to $14\times 14$ patches;
  • Affine base features are equipped with learnable $(x,y)$ per-feature location;
Such model has ~20K parameters, achieves ~98% standard and ~80% adversarial accuracy (AutoAttack) and can still be considered white box as shown in the following figures:
Figure 7: Affine base features learned on full MNIST.
Figure 8: Masked non-negative weights of linear probe layer - per output class.
Figure 9: Sum of affine features weighted by positive weights of linear probe - per output class.